# Empowering LLMs with Pseudo-Untrimmed Videos for Audio-Visual Temporal Understanding

### v2

@@ Introduction
We utilize a text encoder $E_{\text{txt}}$ to embed captions from the video-caption pairs $\{(v_{1},c_{1}),(v_{2},c_{2}),\ldots,(v_{n},c_{n})\}$ sourced from the VALOR-32K dataset. For each caption $c_{i}$, the embedding is given by: $\bar{c}_{i}=E_{\text{txt}}(c_{i})$ Next, we apply a clustering algorithm to the set of embeddings $\{\bar{c}_{1},\bar{c}_{2},\ldots,\bar{c}_{n}\}$ to identify and group videos with similar events, resulting in clusters denoted as $K=\{k_{1},k_{2},\ldots,k_{m}\}$. From each video cluster $k$, we random select one and then identify the $m-1$ top most similar clips within the same cluster. This selection is defined by:
@@
To ensure diverse temporal relationships between the clips, for each selected video $v_{i}\in S_{k}$, we randomly scale its duration within the range $[T_{\text{min}},T_{\text{max}}]$ as
@@
$[t_{\text{start},i}^{\text{new}},~{}t_{\text{end},i}^{\text{new}}]=[T_{\text{offset,i}},~{}T_{\text{offset,i}}+d_{i}\cdot\epsilon]$ (4)
where $T_{\text{offset}}$ is the cumulative duration of all preceding videos in $U_{k}$ after scaling, and the scaling factor $\epsilon=T_{v_{i}}^{\text{new}}/T_{v_{i}}$. We annotate temporal intervals $[t_{\text{start},i}^{\text{new}},t_{\text{end},i}^{\text{new}}]$ within $U_{k}$ to correspond with the content described by the captions $c_{i}$.
@@
$\psi_{t}=\begin{cases}\tilde{a}_{\lceil t/(\omega_{\rho}+1)\rceil}&\text{if}~{}t\mod(\omega_{\rho}+1)\equiv 0,\\ \tilde{f}_{\lceil t/(\omega_{\rho}+1)\rceil}&\text{otherwise},\end{cases}$ (5)
@@
We uniformly extract a minimum of 100 frames from each video to create an interleaved sequence of audio-visual tokens. Video and audio frames are synchronized through interpolation or downsampling to match the Audio-Interleaving Rate, resulting in a sequence of 100 tokens, each representing 1% of the audio-visual video duration. The video scaling factor range is set to $[T_{\text{min}},T_{\text{max}}]=[0.5,2]$ with a step of 0.1, allowing playback speed adjustments from half to double the original rate. Detailed information on clustering and fine-tuning settings, including specific learning rates and epochs, is provided in the Appendix.

### v3

@@ Introduction
We utilize a text encoder $E_{\text{txt}}$ to embed captions from the video-caption pairs $\{(v_{1},c_{1}),(v_{2},c_{2}),\ldots,(v_{n},c_{n})\}$ sourced from the VALOR-32K dataset. For each caption $c_{i}$, the embedding is given by: $\bar{c}_{i}=E_{\text{txt}}(c_{i})$ Next, we apply a clustering algorithm to the set of embeddings $\{\bar{c}_{1},\bar{c}_{2},\ldots,\bar{c}_{n}\}$ to identify and group videos with similar events, resulting in clusters denoted as $K=\{k_{1},k_{2},\ldots,k_{m}\}$. From each video cluster $k$, we random select one and then identify the m1m-1 top most similar clips within the same cluster. This selection is defined by:
@@
To ensure diverse temporal relationships between the clips, for each selected video $v_{i}\in S_{k}$, we randomly scale its duration within the range $T_{\text{min}},T_{\text{max}}]$ as
@@
$t_{\text{start},i}^{\text{new}},~t_{\text{end},i}^{\text{new}}]=[T_{\text{offset,i}},~T_{\text{offset,i}}+d_{i}\cdot\epsilon]$ (4)
where $T_{\text{offset}}$ is the cumulative duration of all preceding videos in $U_{k}$ after scaling, and the scaling factor $\epsilon=T_{v_{i}}^{\text{new}}/T_{v_{i}}$. We annotate temporal intervals $t_{\text{start},i}^{\text{new}},t_{\text{end},i}^{\text{new}}]$ within $U_{k}$ to correspond with the content described by the captions $c_{i}$.
@@
$\psi_{t}=\begin{cases}\tilde{a}_{\lceil t/(\omega_{\rho}+1)\rceil}&\text{if}~t\hskip-7.0pt\mod(\omega_{\rho}+1)\equiv 0,\\ \tilde{f}_{\lceil t/(\omega_{\rho}+1)\rceil}&\text{otherwise},\end{cases}$ (5)
@@
We uniformly extract a minimum of 100 frames from each video to create an interleaved sequence of audio-visual tokens. Video and audio frames are synchronized through interpolation or downsampling to match the Audio-Interleaving Rate, resulting in a sequence of 100 tokens, each representing 1% of the audio-visual video duration. The video scaling factor range is set to $T_{\text{min}},T_{\text{max}}]=[0.5,2]$ with a step of 0.1, allowing playback speed adjustments from half to double the original rate. Detailed information on clustering and fine-tuning settings, including specific learning rates and epochs, is provided in the Appendix.